OpenStudy (josee):

lim as x-> 0+ to (4-(1/sqr.of x))/ (4+(5/sqr.of. x)) um should the first step be to multiply by the congiant?

OpenStudy (amistre64):

multiplying with giants may or maynot be usefu;

OpenStudy (amistre64):

\[\lim_{x->0+} \cfrac{4-\frac{1}{\sqrt{x}}}{4+\cfrac{5}{\sqrt{x}}}\]

OpenStudy (amistre64):

turn it into a normal looking fraction by getting like denoms and putting it into a good looking form

OpenStudy (josee):

Whenever i do a table, it goes to -\[\infty\] until it hits 0 where it doesnt have a numerical value. Would it be that the limit doesnt exist as x approaches 0 from the right?

OpenStudy (amistre64):

\[\lim_{x->0+} \cfrac{\cfrac{4\sqrt{x}-1}{\sqrt{x}}}{\cfrac{4\sqrt{x}+5}{\sqrt{x}}}\]cancel out the like terms: 1/sqrt(x) \[\lim_{x->0+} \cfrac{4\sqrt{x}-1}{4\sqrt{x}+5}\]

OpenStudy (amistre64):

now when x=0; we have a value that we can compare to; namely -1/5